OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

OpenStudy (loser66):

use draw box below to write the problem, please. I don't understand a word

OpenStudy (asnaseer):

I think the asker is asking this:\[(-4x)^0=1\]\[-4x^0=-4\]

OpenStudy (asnaseer):

@Bea_26 -- am I right?

OpenStudy (anonymous):

Yep

OpenStudy (asnaseer):

ok, are you aware that \[a^0=1\]

OpenStudy (anonymous):

Yep

OpenStudy (asnaseer):

for any \(a\) (apart from zero which has some controversy)

OpenStudy (asnaseer):

ok, good, so can you see that:\[(-4x)^0=(a)^0=a^0=1\]if we let \(a=-4x\)

OpenStudy (anonymous):

order of operations

OpenStudy (asnaseer):

@Bea_26 -- Have you heard of PEMDAS

OpenStudy (anonymous):

@asnaseer Yes I Have.

OpenStudy (asnaseer):

that is what @pgpilot326 is trying to explain now

OpenStudy (asnaseer):

in the second case you have:\[-4x^0\]so you need to perform the Exponent operation first, and then the multiplication to get:\[-4x^0=-4\times x^0=-4\times 1=-4\]

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

A Little.

OpenStudy (asnaseer):

which part is troubling you?

OpenStudy (anonymous):

we know that any number raised to the power of zero is one. but the exponent only assume the value right below it so -4x^0 actually means -4* x^0 but when you use parenthesis such as in the case of (-4x)^0 you have to treat -4x as a single value since -4x must be evaluated before evaluating the power. in other words (-4x)^0 = -4^0 * x^0 = 1* 1 = 1

OpenStudy (anonymous):

@asnaseer , The Last Part.

OpenStudy (asnaseer):